# Quantum tunneling from excited states: Recovering imaginary-time instantons from a real-time analysis

### v1

@@ Abstract
We revisit the path integral description of quantum tunneling and show how it can be generalized to excited states. For clarity, we focus on the simple toy model of a point particle in a double-well potential, for which we perform all steps explicitly. Instead of performing the familiar Wick rotation from physical to imaginary time—which is inconsistent with the requisite boundary conditions when treating tunneling from excited states—we regularize the path integral by adding an infinitesimal complex contribution to the Hamiltonian, while keeping time strictly real. We find that this gives rise to a complex stationary-phase solution, in agreement with recent insights from Picard-Lefshitz theory. We then show that there exists a class of analytic solutions for the corresponding equations of motion, which can be made to match the appropriate boundary conditions in the physically relevant limits of a vanishing regulator and an infinite physical time. We provide a detailed discussion of this non-trivial limit. We find that, for systems without an explicit time-dependence, our approach reproduces the picture of an instanton-like solution defined on a finite Euclidean-time interval. Lastly, we discuss the generalization of our approach to broader classes of systems, for which it serves as a reliable framework for high-precision calculations.
@@ I Introduction
A straightforward generalization of this simple picture would suggest that the decay rate out of an excited state should be determined by the Euclidean action of a suitable instanton solution connecting the initial position of interest with the emergence point on the other side of the potential barrier. As the initial state is no longer a local extremum, this motion can be expected to occur within some finite imaginary-time interval, whose duration can be expected to be of the order of the potential’s natural time-scale. In other words, for tunneling out of an excited state, one might expect that the imaginary-time interval of interest would be determined by the form of the potential and the initial state itself. This picture agrees with the result obtained through the WKB approximation, in which the imaginary time appears as a formal tool rather than a representation of the physical time [2, 1, 11, 12].
In the direct approach, however, the imaginary time is usually linked to the real, physical time, from which it arises through a Wick rotation. This would seem to suggest that the length of the Euclidean-time interval is determined by the physical time that has passed since the particle started out from its initial state. Taking seriously this interpretation, our previous observations would suggest that the instanton only exists for one unique (physical) time, of order of the system’s natural time-scale. This conclusion not only appears counterintuitive on its own, but also disagrees with the necessity of a physical time that is significantly longer than the system’s natural time-scale, as found in Refs. [5, 6].
Clearly, the natural candidate for the origin of this issue is the Wick rotation, which usually implicitly relies on the assumption that the system approaches a minimum of the potential at asymptotic physical times. Thus, when interested in the tunneling rate out of an excited state, one is forced to work in real time. To make this manifest, as we demonstrate here, one may instead regularize the path integral by introducing an infinitesimal imaginary part of the Hamiltonian. For time-independent systems, this is equivalent to performing an infinitesimal Wick rotation and eventually taking the appropriate limit. [^1] The interpretation of a complex energy arises naturally when deriving the regularization from the properties of a physical initial state (see, e.g., Refs. [13, 14]). The conceptual subtleties linked to the common step of assigning the imaginary part to the time variable are illustrated, e.g., in Ref. [15]. However, this equivalence is broken for systems with an explicit time-dependence, for which the usual Wick rotation is even more subtle.
@@ II Complex time in quantum tunneling
Eq. (4) implies that the exponent of the decay rate is, to leading order and up to the term $E\cdot\Delta\tau_{I}$, given by the Euclidean action of the solution $\bar{x}_{I}$. The latter can be recognized as the familiar instanton, while the $E\cdot\Delta\tau_{I}$ term can be linked to the fact that the initial and final states are energy eigenstates [42, 43, 44, 45]. For our purpose, the most interesting insight gained from this discussion is that the (finite) length of the Euclidean time interval $\Delta\tau_{I}$ is by construction such that a suitable instanton connecting $x_{i}$ and $x_{s}$ can exist.
Whereas this approach successfully captures the leading-order tunneling rate in many situations, in particular out of the false vacuum, its range of application is limited due to its reliance on the WKB approximation. For example, on a conceptual level, there is no known procedure to extend this approach to incorporate loop corrections or dynamical effects. Even at leading order, the WKB approach breaks down in regions for which $V(x)\sim E$. This includes, in particular, tunneling from an initial state that is relatively close to the top of the potential barrier [11, 12]. This problem is of particular significance for tunneling out of relatively flat minima. Potentials with such features have recently gained interest in the context of new approaches to beyond-the-Standard Model (BSM) model building [46].
@@
When considering tunneling out of the false vacuum, these results are perfectly consistent with the one obtained from the WKB approximation. In particular, they successfully reproduce the infinite imaginary-time interval of the instanton in the limit of large physical times $T\to\infty$ motivated in Refs. [5, 6]. However, it is easy to see that this procedure cannot be used to describe tunneling out of an excited state, for which the Wick rotation would seem to require an infinitely long imaginary-time interval, given the infinite-$T$ limit. Comparing with Eq. (4), we see that taking the limit $\Delta\tau_{I}\rightarrow\infty$ is inconsistent with the WKB approximation.
In addition, performing the usual Wick rotation and taking the limit $\Delta\tau_{I}\rightarrow\infty$ would prevent us from using the saddle-point approximation to evaluate the path integrals in the first place. The initial condition linked to the excited state implies that the instanton begins and ends its motion in points for which $V^{\prime}\neq 0$. This, in turn, implies that the particle’s periodic motion occurs within some finite imaginary time, of order of the system’s dynamical time-scale [42, 43, 44, 45]. In other words, if one were to perform the usual Wick rotation to imaginary time, the tunneling could only be described through the saddle-point approximation for some unique (real) time $T$, and, in particular, not in the physically relevant limit $T\to\infty$. [^2] This issue cannot be avoided by taking into account the possibility of multi-instantons, as the limit of an infinite imaginary time would correspond to an infinite number of instantons, and hence, an infinite Euclidean action.
@@
To understand the origin of this contradiction, we note that, on a qualitative level, the cases of tunneling from the false vacuum and from an excited state differ by the asymptotic behavior of the instanton: whereas in the first case the instanton converges toward the false vacuum, the latter undergoes an infinite, periodic motion, rendering the Wick rotation highly non-trivial. A simple way around this issue would be to avoid the Wick rotation entirely and evaluate the path integral in real time. This idea has been pursued in Ref. [39], which analyzes tunneling out of a false vacuum by performing a Wick rotation $t\to e^{-i\epsilon}t$ and taking the limit $\epsilon\to 0$. For the special case of tunneling out of the false vacuum, and thus considering an infinitely long time interval, these authors find that the decay rate is independent of $\epsilon$.
@@
Most important for our purpose, we observe that the evolution of the amplitude of the real part of the complex-valued instanton reproduces the behavior of the imaginary-time instanton under the identification $\tau=\epsilon t$. In simple terms, this suggests that the transition to a complex-time argument involves a projection rather than a rotation, see Fig. 1. For tunneling out of the false vacuum, this still ensures the convergence of the complex-time solution, as the limit $T\to\infty$ amounts to $\tau\to\infty$ for any non-vanishing $\epsilon$. As we will see, the generalization of this observation will turn out to be crucial for the explanation of the finite imaginary-time support of the instanton corresponding to tunneling out of an excited state.
@@ III Tunneling rate from first principles
We consider a point particle in a potential with a false vacuum $x_{\rm FV}$, surrounded by the false vacuum basin $\Omega_{\rm FV}$. In the most general case, the initial state at $t=0$ of such a particle can be described through a density matrix $\rho$, although we will restrict ourselves to the case of a pure initial state, $\rho=|x_{i}\rangle\langle x_{i}|$. The probability to recover the point particle within some region $\Omega$ (such as $\Omega_{\rm FV}$) at some later time $t=T$ is simply given by [24]
@@
Considering the case $\Omega=\Omega_{\rm FV}$, the probability can be expected to decay as
$P_{\Omega_{\rm FV}}(T)=P_{\Omega_{\rm FV}}(0)\cdot e^{-\Gamma T},$ (8)
$\Gamma(T)=-\frac{1}{P_{\Omega_{\rm FV}}(T)}\frac{\text{d}}{\text{d}T}P_{\Omega_{\rm FV}}(T)=\frac{1}{P_{\Omega_{\rm FV}}(T)}\frac{\text{d}}{\text{d}t}P_{R}(T),$ (9)
where $R$ is the region in position space into which the particle tunnels. Henceforth we adopt the abbreviation $Z(T)\equiv P_{\Omega_{\rm FV}}(T)$. The last of these factors can be represented through the point particle’s propagator $D_{F}(x_{i},t_{i}|x_{f},t_{f})$ as
$P_{R}=$ $\int_{R}{\rm d}x_{f}\ D_{F}(x_{f},T|x_{i})D_{F}(x_{i}|x_{f},T).$ (10)
@@
where $x_{s}$ is the closest point in $R$ for which the particle’s energy is degenerate with its energy at point $x_{i}$. [^3] A similar decomposition is possible in the presence of more degrees of freedom, in which case one needs to include an additional integral of $x_{s}$ over a suitable hypersurface. To define the auxiliary quantity $\bar{D}_{F}$, we may first introduce the functional $F_{x_{s}}[x]$, which (as noted above) maps any time-dependent path $x$ onto the time when it first reaches $x_{s}\subset R$. In terms of this object, $\bar{D}_{F}$ is defined as
$\bar{D}_{F}(x_{i}|x_{s},t_{s})\equiv\mathcal{N}\underset{x(0)=x_{i}}{\overset{x(t_{s})=x_{s}}{\int}}{\mathcal{D}}x\ e^{iS[x]}\delta\left(F_{x_{s}}[x]-t_{s}\right),$ (12)
@@
$P_{R}=$ $\int_{0}^{T}\text{d}t_{s}\int_{0}^{T}\text{d}t_{s}^{\prime}\,\bar{D}_{F}(x_{i}|x_{s},t_{s})\bar{D}_{F}^{*}(x_{i}|x_{s},t_{s}^{\prime})$
@@
$\Gamma=$ $\lim_{T\to\infty}\frac{|\mathcal{N}|^{2}}{Z(T)}\underset{x_{1}(0)=x_{i}}{\overset{x_{1}(T)=x_{s}}{\int}}{\mathcal{D}}x_{1}\ e^{iS[x_{1}]}\delta\left(F_{x_{s}}[x_{1}]-T\right)$
×$\quad\quad\quad\quad\quad\quad\times\left(\ \ \ \underset{x_{2}(0)=x_{i}}{\overset{x_{2}(T)=x_{s}}{\int}}{\mathcal{D}}x_{2}\ e^{iS[x_{2}]}\right)^{*}+c.c.$ (17)
@@
$Z(T)=$ $P_{\Omega_{\rm FV}}(T)=\int_{\Omega_{\rm FV}}{\rm d}x_{v}\langle x_{v},T|x_{i}\rangle\langle x_{i}|x_{v},T\rangle$
$=$ $\int_{\Omega_{\rm FV}}{\rm d}x_{v}\ D_{F}(x_{v},T|x_{i})D_{F}(x_{i}|x_{v},T)$
$=$ $|\mathcal{N}|^{2}\int_{\Omega_{\rm FV}}{\rm d}x_{v}\underset{z_{1}(0)=x_{i}}{\overset{z_{1}(T)=x_{v}}{\int}}{\mathcal{D}}z_{1}\ e^{iS[z_{1}]}\left(\ \ \ \underset{z_{2}(0)=x_{i}}{\overset{z_{2}(T)=x_{v}}{\int}}{\mathcal{D}}z_{2}\ e^{iS[z_{2}]}\right)^{*}.$ (18)
In the cases $x_{i}=x_{\rm FV}$ and $z_{i}=x_{\rm FV}$, these expressions can be simplified through Wick rotations, which translate the real-time actions to Euclidean ones. Crucially, this suggests that one identify the physical time $T$ with the duration of the imaginary-time interval $\tau$. Motivated by our previous discussion, we will instead evaluate these path integrals through a saddle-point approximation in real time.
In Secs. V and  VI we show how to construct suitable stationary-phase solutions for the equations of motions, with the boundary conditions imposed by the path integral. However, even without having at hand the precise form of these solutions, it is easy to see that they are fully determined by their boundary conditions. Formally, this implies $\bar{x}_{1}=\bar{x}_{2}\equiv\bar{x}$, where $\bar{x}_{1,2}$ is the stationary phase of the $x_{1,2}$ path integral. As the same holds true for the path integrals in the denominator involving paths $z_{1,2}$ with corresponding stationary phases $\bar{z}_{1,2}$ (restricted to paths that terminate within $\Omega_{\rm FV}$), we thus find that the decay rate is, to leading, of the form

### v2

# Toward quantum tunneling from excited states: Recovering imaginary-time instantons from a real-time analysis

@@ Abstract
We revisit the path integral description of quantum tunneling and lay the groundwork for its generalization to excites states through real-time path integral techniques. For clarity, we focus on the simple toy model of a point particle in a double-well potential, for which we perform all steps explicitly. Instead of performing the familiar Wick rotation from physical to imaginary time—which is inconsistent with the requisite boundary conditions when treating tunneling from states other than the false vacuum—we regularize the path integral by adding an infinitesimal complex contribution to the Hamiltonian, while keeping time strictly real. We find that this gives rise to a complex stationary-phase solution, in agreement with recent insights from Picard-Lefshitz theory. We then show that there exists a class of analytic solutions for the corresponding equations of motion, which can be made to match the appropriate boundary conditions in the physically relevant limits of a vanishing regulator and an infinite physical time. We provide a detailed discussion of this non-trivial limit. We find that, for systems without an explicit time-dependence, our approach reproduces the picture of an instanton-like solution defined on a finite Euclidean-time interval. Lastly, we discuss the generalization of our approach to broader classes of systems, for which it serves as a reliable framework for high-precision calculations.
@@ I Introduction
A straightforward generalization of this simple picture would suggest that the decay rate out of an excited state should be determined by the Euclidean action of a suitable instanton solution connecting the initial position of interest with the emergence point on the other side of the potential barrier. As the initial state is no longer a local extremum, this motion can be expected to occur within some finite imaginary-time interval, whose duration should be of the order of the potential’s natural time-scale. In other words, for tunneling out of a state different from the false vacuum, one might expect that the imaginary-time interval of interest would be determined by the form of the potential and the initial state itself. This picture agrees with the result obtained through the WKB approximation, in which the imaginary time appears as a formal tool rather than a representation of the physical time [2, 1, 11, 12].
In the direct approach, however, the imaginary time is usually linked to the real, physical time, from which it arises through a Wick rotation. This would seem to suggest that the length of the Euclidean-time interval is determined by the physical time that has passed since the particle started out from its initial state. Taking seriously this interpretation, our previous observations would then imply that the instanton only exists for one unique (physical) time, of order of the system’s natural time-scale. This conclusion not only appears counterintuitive on its own, but also disagrees with the necessity of a physical time that is significantly longer than the system’s natural time-scale, as found in Refs. [5, 6].
Clearly, the natural candidate for the origin of this issue is the Wick rotation, which usually implicitly relies on the assumption that the system approaches a minimum of the potential at asymptotic physical times. Thus, when interested in the tunneling rate out of a general state, one is forced to work in real time. To make this manifest, as we demonstrate here, one may instead regularize the path integral by introducing an infinitesimal imaginary part of the Hamiltonian. For time-independent systems, this is equivalent to performing an infinitesimal Wick rotation and eventually taking the appropriate limit. [^1] The interpretation of a complex energy arises naturally when deriving the regularization from the properties of a physical initial state (see, e.g., Refs. [13, 14]). The conceptual subtleties linked to the common step of assigning the imaginary part to the time variable are illustrated, e.g., in Ref. [15]. However, this equivalence is broken for systems with an explicit time-dependence, for which the usual Wick rotation is even more subtle.
@@ II Complex time in quantum tunneling
Eq. (4) implies that the exponent of the decay rate is, to leading order and up to the term $E\cdot\Delta\tau_{I}$, given by the Euclidean action of the solution $\bar{x}_{I}$. The latter can be recognized as the familiar instanton, while the $E\cdot\Delta\tau_{I}$ term can be linked to the fact that the initial and final states are resonance states [42, 43, 44, 45]. For our purpose, the most interesting insight gained from this discussion is that the (finite) length of the Euclidean time interval $\Delta\tau_{I}$ is by construction such that a suitable instanton connecting $x_{i}$ and $x_{s}$ can exist.
Whereas this approach successfully captures the leading-order tunneling rate in many situations, in particular out of the false vacuum, its range of application is limited due to its reliance on the WKB approximation. For example, on a conceptual level, there is no known procedure to extend this approach to incorporate loop corrections or dynamical effects. Even at leading order, the WKB approach breaks down in regions for which $V(x)\sim E$. This includes, in particular, tunneling from an initial state that is relatively close to the top of the potential barrier [11, 12]. This problem is of particular significance for tunneling out of relatively flat minima. Potentials with such features have recently gained interest in the context of new approaches to beyond-the-Standard Model (BSM) model building [46, 47].
@@
[…]
In addition, performing the usual Wick rotation and taking the limit $\Delta\tau_{I}\rightarrow\infty$ would prevent us from using the saddle-point approximation to evaluate the path integrals in the first place. The initial condition implies that the instanton begins and ends its motion in points for which $V^{\prime}\neq 0$. This, in turn, implies that the particle’s periodic motion occurs within some finite imaginary time, of order of the system’s dynamical time-scale [42, 43, 44, 45]. In other words, if one were to perform the usual Wick rotation to imaginary time, the tunneling could only be described through the saddle-point approximation for some unique (real) time $T$, and, in particular, not in the physically relevant limit $T\to\infty$. [^2] This issue cannot be avoided by taking into account the possibility of multi-instantons, as the limit of an infinite imaginary time would correspond to an infinite number of instantons, and hence, an infinite Euclidean action.
@@
To understand the origin of this contradiction, we note that, on a qualitative level, the case of tunneling from the false vacuum is unique due to the asymptotic behavior of the corresponding instanton: it converges towards the false vacuum, whereas a generic solution would undergo an infinite, periodic motion. A simple way around the complications arising from the latter case would be to avoid the Wick rotation entirely and evaluate the path integral in real time. This idea has been pursued in Ref. [39], which analyzes tunneling out of a false vacuum by performing a Wick rotation $t\to e^{-i\epsilon}t$ and taking the limit $\epsilon\to 0$. For the special case of tunneling out of the false vacuum, and thus considering an infinitely long time interval, these authors find that the decay rate is independent of $\epsilon$.
@@
Most important for our purpose, we observe that the evolution of the amplitude of the real part of the complex-valued instanton reproduces the behavior of the imaginary-time instanton under the identification $\tau=\epsilon t$. In simple terms, this suggests that the transition to a complex-time argument involves a projection rather than a rotation, see Fig. 1. For tunneling out of the false vacuum, this still ensures the convergence of the complex-time solution, as the limit $T\to\infty$ amounts to $\tau\to\infty$ for any non-vanishing $\epsilon$. As we will see, the generalization of this observation will turn out to be crucial for the explanation of the finite imaginary-time support of the instanton corresponding to tunneling out of a state other than the false vacuum.
@@ III Tunneling rate from first principles
III.1 Defining the tunneling rate
We consider a point particle in a potential with two vacua $x_{\pm}$, surrounded by two basins $\Omega_{\pm}$. See Fig. 2. In the most general case, the initial state at $t=0$ of such a particle can be described through a density matrix $\rho=\int{\rm d}x_{i}\int{\rm d}x_{j}\ \rho_{ij}|x_{i}\rangle\langle x_{j}|$. For a pure state with wave function $\psi(x)$, the coefficient $\rho_{ij}$ takes the simple form $\rho_{ij}=\psi(x_{i})\psi^{*}(x_{j})$. For the remainder of this article, we will restrict ourselves to the idealized case of a highly localized (but on its own unphysical) initial state, $\rho=|x_{i}\rangle\langle x_{i}|$. For our present analysis, we consider our simplified choice because it allows for a clear understanding of many conceptual aspects of our computation. Moreover, any density matrix can be expressed in terms of matrix elements $|x_{i}\rangle\langle x_{j}|$. As argued in Ref. [14], symmetry suggests that for generic systems the dominant contribution will arise from stationary phase solutions with $x_{i}=x_{j}$. Whenever this is the case, this allows to represent the tunneling rate out of any initial state as the superposition of the rates out of individual position eigenstates, each of which can be calculated using our results—in particular, excited states. However, even in cases in which non-diagonal elements of $\rho_{ij}$ should become important we anticipate that it is straightforward to generalize our results accordingly. Crucially, the state we consider is of course itself not a physical state, but rather serves the purpose of making transparent the semi-classical structure of the relevant class of path integrals. For an example of an actual physical state, consider Ref. [14]
Figure 2: A visualization of the ansatz of Eq. (7). At $t=0$, the point particle is localized at $x_{i}$ (black curve). Starting from this position, it will either tunnel to region $\Omega_{+}$ with an exponentially suppressed probability (solid blue peak in region $\Omega_{+})$, or move to the left. Following the latter path, the particle will reach a turning point (solid red peak within $\Omega_{-}$) after some time $\Delta t_{R}/2$. The particle would then oscillate back toward $x_{i}$, returning with some diminished probability (green curve). Then the process will repeat, with correspondingly smaller probabilities both to tunnel into $\Omega_{+}$ (dashed blue curve) or to oscillate to the left within $\Omega_{-}$ (dashed red curve).
We may now anticipate the relevant behavior of the system using the interpretation of a path integral as a superposition of all possible paths, dominated by the particle’s classical trajectory. First, we consider a particle at rest at the position $x_{i}$ at $t=0$. Classically, such a particle can then be expected to perform an oscillatory motion within the basin $\Omega_{-}$. We can furthermore expect that it is most likely for this particle to tunnel into the basin $\Omega_{+}$ at the times when it is close to the turning point $x_{i}$. During these times, the total probability transfer can be expected to be proportional to the total probability within $\Omega_{-}$, $\Delta P_{\Omega_{-}}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-}\mathcal{T}\cdot P_{\Omega_{-}}$, with the usual transmission coefficient $\mathcal{T}$ usually calculated through the WKB method. Thus, during these time intervals, we expect the probability to find the particle within the basin $\Omega_{-}$ to evolve with time as
$P_{\Omega_{-}}(T)=P_{\Omega_{-}}(0)\cdot e^{-\Gamma T},$ (7)
$\Gamma(T)=-\frac{1}{P_{\Omega_{-}}(T)}\frac{\text{d}}{\text{d}T}P_{\Omega_{-}}(T)=\frac{1}{P_{\Omega_{-}}(T)}\frac{\text{d}}{\text{d}t}P_{\Omega_{+}}(T),$ (8)
where $\Omega_{+}$ is the region in position space into which the particle tunnels. Henceforth we adopt the abbreviation $Z(T)\equiv P_{\Omega_{-}}(T)$. For the remainder of this article, we will use the quantity $\Gamma$ as the definition of the tunneling rate.
The probability to recover the point particle within some region $\Omega$ (such as $\Omega_{-}$) at some later time $t=T$ is simply given by [24]
@@ III.1 Defining the tunneling rate
An important subtlety of this ansatz is that we have arrived at it by considering a path in which the particle starts with vanishing initial momentum. It is, however, straightforward to see that the above behavior can be found for any initial momentum, albeit with different times $\Delta t_{R}$. Conversely, for any given time $T$, the loss of probability within $\Omega_{-}$ is of the form of Eq. (7), where $\Gamma$ is determined by the contribution representing a different initial momentum. We will revisit this matter in Sec. V.2, where we will show how this picture emerges in a natural way from our calculation.
III.2 Path Integral representation
The last factor in Eq. (8) can be represented through the point particle’s propagator $D_{F}(x_{i},t_{i}|x_{f},t_{f})$ as
$P_{\Omega_{+}}=$ $\int_{\Omega_{+}}{\rm d}x_{f}\ D_{F}(x_{f},T|x_{i})D_{F}(x_{i}|x_{f},T).$ (10)
@@ III.2 Path Integral representation
where $x_{s}$ is the closest point in $\Omega_{+}$ for which the particle’s energy is degenerate with its energy at point $x_{i}$. [^3] A similar decomposition is possible in the presence of more degrees of freedom, in which case one needs to include an additional integral of $x_{s}$ over a suitable hypersurface. To define the auxiliary quantity $\bar{D}_{F}$, we may first introduce the functional $F_{x_{s}}[x]$, which (as noted above) maps any time-dependent path $x$ onto the time when it first reaches $x_{s}\subset\Omega_{+}$. In terms of this object, $\bar{D}_{F}$ is defined as
$\bar{D}_{F}(x_{i}|x_{s},t_{s})\equiv\mathcal{N}\hskip-10.00002pt\underset{x(0)=x_{i}}{\overset{x(t_{s})=x_{s}}{\int}}\hskip-10.00002pt{\mathcal{D}}x\ e^{iS[x]}\delta\left(F_{x_{s}}[x]-t_{s}\right),$ (12)
@@
$P_{\Omega_{+}}=$ $\int_{0}^{T}\hskip-3.99994pt\text{d}t_{s}\int_{0}^{T}\hskip-3.99994pt\text{d}t_{s}^{\prime}\,\bar{D}_{F}(x_{i}|x_{s},t_{s})\bar{D}_{F}^{*}(x_{i}|x_{s},t_{s}^{\prime})$
@@
$\Gamma=$ $\lim_{T\to\infty}\frac{|\mathcal{N}|^{2}}{Z(T)}\hskip 0.0pt\underset{x_{1}(0)=x_{i}}{\overset{x_{1}(T)=x_{s}}{\int}}\hskip 0.0pt{\mathcal{D}}x_{1}\ e^{iS[x_{1}]}\delta\left(F_{x_{s}}[x_{1}]-T\right)$
×$\quad\quad\quad\quad\quad\quad\times\left(\ \ \ \hskip-10.00002pt\underset{x_{2}(0)=x_{i}}{\overset{x_{2}(T)=x_{s}}{\int}}\hskip-10.00002pt{\mathcal{D}}x_{2}\ e^{iS[x_{2}]}\right)^{*}+c.c.$ (17)
@@
$Z(T)=$ $P_{\Omega_{-}}(T)=\int_{\Omega_{-}}{\rm d}x_{v}\langle x_{v},T|x_{i}\rangle\langle x_{i}|x_{v},T\rangle$ (18)
$=$ $\int_{\Omega_{-}}{\rm d}x_{v}\ D_{F}(x_{v},T|x_{i})D_{F}(x_{i}|x_{v},T)$
$=$ $|\mathcal{N}|^{2}\int_{\Omega_{-}}{\rm d}x_{v}\hskip-10.00002pt\underset{z_{1}(0)=x_{i}}{\overset{z_{1}(T)=x_{v}}{\int}}\hskip-10.00002pt{\mathcal{D}}z_{1}\ e^{iS[z_{1}]}\left(\ \ \ \hskip-10.00002pt\underset{z_{2}(0)=x_{i}}{\overset{z_{2}(T)=x_{v}}{\int}}\hskip-10.00002pt{\mathcal{D}}z_{2}\ e^{iS[z_{2}]}\right)^{*}.$
In the cases $x_{i}=x_{-}$ and $z_{i}=x_{-}$, these expressions can be simplified through Wick rotations, which translate the real-time actions to Euclidean ones. Crucially, this suggests that one identify the physical time $T$ with the duration of the imaginary-time interval $\tau$. Motivated by our previous discussion, we will instead evaluate these path integrals through a saddle-point approximation in real time.
In Secs. V and  VI we show how to construct suitable stationary-phase solutions for the equations of motions, with the boundary conditions imposed by the path integral. However, even without having at hand the precise form of these solutions, it is easy to see that they are fully determined by their boundary conditions. Formally, this implies $\bar{x}_{1}=\bar{x}_{2}\equiv\bar{x}$, where $\bar{x}_{1,2}$ is the stationary phase of the $x_{1,2}$ path integral. As the same holds true for the path integrals in the denominator involving paths $z_{1,2}$ with corresponding stationary phases $\bar{z}_{1,2}$ (restricted to paths that terminate within $\Omega_{-}$), we thus find that the decay rate is, to leading, of the form